THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 106 — 2 August 2010

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Editorial

Dear Colleagues,

It is my pleasure to present you the 106^{th} issue of the Magellanic Clouds Newsletter. Besides five papers on periodluminosity relations and five on supernova remnants, some on 30 Doradus and (its) massive (runaway) stars, and lots of other very interesting results, there are some quite original works this time, e.g., reporting diffuse far-UV emission, radio emission from dark matter annihilation, the Magellanic Stream feeding the Halo with hot plasma, and a novel technique of ”carbon dating” B[e] stars.

The University of Hertfordshire offers postdoctoral fellowships, and Maria-Rosa Cioni is especially keen on someone to work with her on the VISTA survey of the Magellanic System. The first data look genuinely impressive!

There is going to be a nice Magellanic Clouds meeting as part of the diverse and exciting conference ”Zooming in: The Cosmos at High Resolution”, in Bonn in September. We are looking forward to seeing many of you there.

The next issue is planned to be distributed on the 1^{st} of October 2010. Front cover pictures remain welcome!

Editorially Yours, Jacco van Loon

Refereed Journal Papers

Cold dust in three massive evolved stars in the LMC

M.L. Boyer^{1}, B. Sargent^{1}, J.Th. van Loon^{2}, S. Srinivasan^{3}, G.C. Clayton^{4}, F. Kemper^{5}, L.J. Smith^{1}, M. Matsuura^{6,7}, Paul M. Woods^{5}, M. Marengo^{8}, M. Meixner^{1}, C. Engelbracht^{9}, K.D. Gordon^{1}, S. Hony^{10}, R.

Indebetouw^{11}, K. Misselt^{9}, K. Okumura^{10}, P. Panuzzo^{10}, D. Riebel^{12}, J. Roman-Duval^{1}, M. Sauvage^{10} and G.C. Sloan^{13}

1Space Telescope Science Institute, USA 2School of Physical and Geographical Sciences, Lennard-Jones Laboratories, Keele University, UK 3Institut d’Astrophysique de Paris, France 4Louisiana State University, Department of Physics and Astronomy, USA 5Jodrell Bank Centre for Astrophysics, Alan Turing Building, School of Physics and Astronomy, University of Manchester, UK 6Department of Physics and Astronomy, University College London, UK 7Mullard Space Science Laboratory, University College London, UK 8Department of Physics and Astronomy, Iowa State University, USA 9Steward Observatory, University of Arizona, USA 10CEA, Laboratoire AIM, Irfu/SAp, France 11National Radio Astronomy Observatory, Department of Astronomy, University of Virginia, USA 12Johns Hopkins University, Department of Physics and Astronomy, USA 13Department of Astronomy, Cornell University, USA

Massive evolved stars can produce large amounts of dust, and far-infrared (IR) data are essential for determining the contribution of cold dust to the total dust mass. Using Herschel, we search for cold dust in three very dusty massive evolved stars in the Large Magellanic Cloud: R 71 is a Luminous Blue Variable, HD 36402 is a Wolf-Rayet triple system, and IRAS 05280−6910 is a red supergiant. We model the spectral energy distributions using radiative transfer codes and find that these three stars have mass-loss rates up to 10^{−3} M_{⊙} yr^{−1}, suggesting that high-mass stars are important contributors to the life-cycle of dust. We found far-IR excesses in two objects, but these excesses appear to be associated with ISM and star-forming regions. Cold dust (T < 100 K) may thus not be an important contributor to the dust masses of evolved stars.

Accepted for publication in A&A Letters (Herschel Special Issue) Available from arXiv:1005.5167

On the formation of multiple stellar populations in globular clusters

Charlie Conroy^{1} and David N. Spergel^{1}

1Princeton University, USA

Nearly all globular clusters (GCs) studied to date show evidence for multiple stellar populations, in stark contrast to the conventional view that GCs are a mono-metallic, coeval population of stars. This generic feature must therefore emerge naturally within massive star cluster formation. Building on earlier work, we propose a simple physical model for the early evolution (several 10^{8} yr) of GCs that is consistent with all of the available data. We consider the effects of stellar mass-loss, type II and prompt type Ia supernovae, ram pressure stripping, and accretion from the ambient interstellar medium (ISM) on the development of a young GC’s own gas reservoir. In our model, type II SNe from a first generation of star formation clears the GC of its initial gas reservoir. Over the next several 10^{8} yr, mass lost from AGB stars and matter accreted from the ambient ISM collect at the center of the GC; this material must remain quite cool (T ∼ 10^{2} K), but does not catastrophically cool on a crossing time because of the high Lyman–Werner flux density in young GCs. The collection of gas within the GC must compete with ram pressure stripping from the ambient ISM. After several 10^{8} yr, the Lyman–Werner photon flux density drops by more than three orders of magnitude, allowing molecular hydrogen and then stars to form. After this second generation of star formation, type II SNe from the second generation and then prompt type Ia SNe associated with the first generation maintain a gas-free GC, thereby ending the cycle of star formation events. Our model makes clear predictions for the presence or absence of multiple stellar populations within GCs as a function of GC mass and environment, and the presence or lack of internal dispersion of Fe-peak and α-elements as a function of cluster age. Analyzing intermediate-age LMC clusters, we find for the first time evidence for a mass threshold of ∼ 10^{4} M_{⊙} below which LMC clusters appear to be truly coeval. This threshold mass is consistent with our predictions for the mass at which ram pressure stripping is capable of clearing gas from clusters in the LMC at the present epoch. Recently, claims have been made that multiple populations within GCs require that GCs form at the center of their own dark matter halos. We argue that such a scenario is implausible, except perhaps for the most massive Galactic GCs (including ω Cen, M 22, M 54, and NGC 2808). Observations of the young and intermediate-age clusters in the LMC and M 31 will provide strong constraints on our proposed scenario.

Submitted to Astrophysical Journal Available from arXiv:1005.4934

On the interpretation of supernova light echo profiles and spectra

A. Rest^{1}, B. Sinnott^{2}, D.L. Welch^{2}, R.J. Foley^{3}, G. Narayan^{1} and K. Mandel^{3}

1Department of Physics, Harvard University, 17 Oxford Street, Cambridge, MA 02138, USA 2Dept. of Physics and Astronomy, McMaster University, Hamilton, Ontario, L8S 4M1, Canada 3Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, USA

The light echo systems of historical supernovae in the Milky Way and local group galaxies provide an unprecedented opportunity to reveal the effects of asymmetry on observables, particularly optical spectra. Scattering dust at different locations on the light echo ellipsoid witnesses the supernova from different perspectives and the light consequently scattered towards Earth preserves the shape of line profile variations introduced by asymmetries in the supernova photosphere. However, the interpretation of supernova light echo spectra to date has not involved a detailed consideration of the effects of outburst duration and geometrical scattering modifications due to finite scattering dust filament dimension, inclination, and image point-spread function and spectrograph slit width. In this paper, we explore the implications of these factors and present a framework for future resolved supernova light echo spectra interpretation, and test it against Cas A and SN 1987A light echo spectra. We conclude that the full modeling of the dimensions and orientation of the scattering dust using the observed light echoes at two or more epochs is critical for the correct interpretation of light echo spectra. Indeed, without doing so one might falsely conclude that differences exist when none are actually present.

Submitted to ApJ Available from arXiv:1004.3783

Massive runaway stars in the Large Magellanic Cloud

V.V. Gvaramadze^{1,2,3}, P. Kroupa^{1} and J. Pflamm-Altenburg^{1}

1Argelander-Institut für Astronomie, Universität Bonn, Auf dem Hügel 73, 53121 Bonn, Germany 2Sternberg Astronomical Institute, Moscow State University, Universitetskij Pr. 13, Moscow 119992, Russia 3Isaac Newton Institute of Chile, Moscow Branch, Universitetskij Pr. 13, Moscow 119992, Russia

The origin of massive field stars in the Large Magellanic Cloud (LMC) has long been an enigma. The recent measurements of large offsets (∼ 100 km s^{−1}) between the heliocentric radial velocities of some very massive (O2-type) field stars and the systemic LMC velocity provides a possible explanation of this enigma and suggests that the field stars are runaway stars ejected from their birth places at the very beginning of their parent cluster’s dynamical evolution. A straightforward way to prove this explanation is to measure the proper motions of the field stars and to show that they are moving away from one of the nearby star clusters or OB associations. This approach however is complicated by the large distance to the LMC, which makes accurate proper motion measurements difficult. We use an alternative approach for solving the problem (first applied for Galactic field stars), based on the search for bow shocks produced by runaway stars. The geometry of detected bow shocks would allow us to infer the direction of stellar motion and thereby to determine their possible parent clusters. In this paper we present the results of a search for bow shocks around six massive field stars which were suggested in the literature as candidate runaway stars. Using archival Spitzer Space Telescope data, we found a bow shock associated with one of our program stars, the O2 V((f*)) star BI 237, which is the first-ever detection of bow shocks in the LMC. Orientation of the bow shock suggests that BI 237 was ejected from the OB association LH 82 (located at ≃ 120 pc in projection from the star). A by-product of our search is the detection of bow shocks generated by four OB stars in the field of the LMC and an arc-like structure attached to the candidate luminous blue variable R 81 (HD 269128). The geometry of two of these bow shocks is consistent with the possibility that their associated stars were ejected from the 30 Doradus star forming complex. We discuss implications of our findings for the problem of the origin of runaway stars and the early dynamical evolution of star clusters.

Accepted for publication in A&A Available from arXiv:1006.0225

Cepheid period–luminosity relation from the AKARI observations

Chow-Choong Ngeow^{1}, Yoshifusa Ita^{2}, Shashi M. Kanbur^{3}, Hilding Neilson^{4}, Takashi Onaka^{5} and Daisuke Kato^{5}

1National Central University, Taiwan 2National Astronomical Observatory of Japan, Japan 3SUNY Oswego, USA 4University of Toronto, Canada 5The University of Tokyo, Japan

In this paper, we derive the period–luminosity (P–L) relation for Large Magellanic Cloud (LMC) Cepheids based on mid-infrared AKARI observations. AKARI’s IRC sources were matched to the OGLE-III LMC Cepheid catalog. Together with the available I band light curves from the OGLE-III catalog, potential false matches were removed from the sample. This procedure excluded most of the sources in the S7 and S11 bands: hence only the P–L relation in the N3 band was derived in this paper. Random-phase corrections were included in deriving the P–L relation for the single epoch AKARI data, even though the derived P–L relation is consistent with the P–L relation without random-phase correction, though there is a ∼ 7 per-cent improvement in the dispersion of the P–L relation. The final adopted N3 band P–L relation is N3 = −3.246log(P) + 15.844, with a dispersion of 0.149.

Accepted for publication in MNRAS Available from arXiv:1006.1553

Period–colour and amplitude–colour relations in classical Cepheid variables — VI. New challenges for pulsation models

Shashi M. Kanbur^{1}, M. Marconi^{2}, Chow-Choong Ngeow^{3}, I. Musella^{2}, M. Turner^{4}, A. James^{5}, S. Magin^{1} and J. Halsey^{1}

1State University of New York at Oswego, USA 2Osservatorio Astronomico di Capodimonte, Italy 3National Central University, Taiwan 4Rice University, USA 5State University of New York at Geneseo, USA

We present multiphase period–color/amplitude–color/period–luminosity relations using OGLE III and Galactic Cepheid data and compare with state of the art theoretical pulsation models. Using this new way to compare models and observations, we find convincing evidence that both period–color and period–luminosity relations as a function of phase are dynamic and highly nonlinear at certain pulsation phases. We extend this to a multiphase Wesenheit function and find the same result. Hence our results cannot be due to reddening errors. We present statistical tests and the urls of movies depiciting the period–color/period–luminosity and Wesenheit relations as a function of phase for the LMC OGLE III Cepheid data: these tests and movies clearly demonstrate nonlinearity as a function of phase and offer a new window towards a deeper understanding of stellar pulsation. When comparing with models, we find that the models also predict this nonlinearity in both period–color and period–luminosity planes. The models with (Z = 0.004, Y = 0.25) fare better in mimicking the LMC Cepheid relations, particularly at longer periods, though the models predict systematically higher amplitudes than the observations.

Accepted for publication in MNRAS Available from arXiv:1006.1821

Supernova remnants and star formation in the Large Magellanic Cloud

K. M. Desai^{1}, Y.-H. Chu^{1}, R. A. Gruendl^{1}, W. Dluger^{1}, M. Katz^{1}, T. Wong^{1}, C.-H.R. Chen^{2}, L.W. Looney^{1}, A. Hughes^{3,4}, E. Muller^{5}, J. Ott^{6} and J.L. Pineda^{7}

1Department of Astronomy, University of Illinois at Urbana-Champaign, USA 2Department of Astronomy, University of Virginia, Charlottesville, USA 3Centre for Supercomputing and Astrophysics, Swinburne University of Technology, Australia 4CSIRO Australia Telescope National Facility, Australia 5Department of Physics and Astrophysics, Nagoya University, Japan 6National Radio Astronomy Observatory, USA 7NASA/JPL, USA

It has often been suggested that supernova remnants (SNRs) can trigger star formation. To investigate the relationship between SNRs and star formation, we have examined the known sample of 45 SNRs in the Large Magellanic Cloud to search for associated young stellar objects (YSOs) and molecular clouds. We find seven SNRs associated with both YSOs and molecular clouds, three SNRs associated with YSOs but not molecular clouds, and eight SNRs near molecular clouds but not associated with YSOs. Among the 10 SNRs associated with YSOs, the association between the YSOs and SNRs can be either rejected or cannot be convincingly established for eight cases. Only two SNRs have YSOs closely aligned along their rims; however, the time elapsed since the SNR began to interact with the YSOs’ natal clouds is much shorter than the contraction timescales of the YSOs, and thus we do not see any evidence of SNR-triggered star formation in the LMC. The 15 SNRs that are near molecular clouds may trigger star formation in the future when the SNR shocks have slowed down to < 45 km s^{−1}. We discuss how SNRs can alter the physical properties and abundances of YSOs.

Accepted for publication in The Astronomical Journal Available from arXiv:1006.3344

Towards an explanation for the 30 Dor (LMC) Honeycomb nebula — the impact of recent observations and spectral analysis

John Meaburn^{1}, Matt Redman^{2}, Panayotis Boumis^{3} and Eamonn Harvey^{2}

1Jodrell Bank Centre for Astrophysics, University of Manchester, UK 2Centre for Astronomy,School of Physics, National University of Ireland Galway, Ireland 3Institute of Astronomy & Astrophysics, National Observatory of Athens, Greece

The unique Honeycomb nebula, most likely a peculiar supernova remnant, lies in 30 Doradus in the Large Magellanic Cloud. Due to its proximity to SN 1987A, it has been serendipitously and intentionally observed at many wavelengths.

Here, an optical spectral analysis of forbidden line ratios is performed in order to compare the Honeycomb high-speed gas with supernova remnants in the Galaxy and the LMC, with galactic Wolf–Rayet nebulae and with the optical line emission from the interaction zone of the SS 433 microquasar and W 50 supernova remnant system. An empirical spatiokinematic model of the images and spectra for the Honeycomb reveals that its striking appearance is most likely due to a fortuitous viewing angle. The Honeycomb nebula is more extended in soft X-ray emission and could in fact be a small part of the edge of a giant LMC shell revealed for the first time in this short wavelength domain. It is also suggested that a previously unnoticed region of optical emission may in fact be an extension of the Honeycomb around the edge of this giant shell. A secondary supernova explosion in the edge of a giant shell is considered for the creation of the Honeycomb nebula. A microquasar origin of the Honeycomb nebula as opposed to a simple supernova origin is also evaluated.

Accepted for publication in MNRAS Available from arXiv:1006.2692

A two-component power law covering nearly four Orders of magnitude in the power spectrum of Spitzer far-infrared emission from the Large

Magellanic Cloud

David L. Block^{1}, Ivânio Puerari^{2}, Bruce G. Elmegreen^{3} and Frédéric Bournaud^{4}

1School of Computational and Applied Mathematics, University of the Witwatersrand, Private Bag 3, WITS 2050, South Africa 2Instituto Nacional de Astrofísica, Optica y Electrónica, Calle Luis Enrique Erro 1, 72840 Santa María Tonantzintla, Puebla, México 3IBM T.J. Watson Research Center, 1101 Kitchawan Road, Yorktown Heights, New York 10598 USA 4CEA, IRFU, SAp, F-91191 Gif-sur-Yvette, France

Power spectra of Large Magellanic Cloud (LMC) emission at 24, 70 and 160 μm observed with the Spitzer Space Telescope have a two-component power-law structure with a shallow slope of −1.6 at low wavenumber, k, and a steep slope of −2.9 at high k. The break occurs at k^{−1} ∼ 100–200 pc, which is interpreted as the line-of-sight thickness of the LMC disk. The slopes are slightly steeper for longer wavelengths, suggesting the cooler dust emission is smoother than the hot emission. The power spectrum covers ∼ 3.5 orders of magnitude and the break in the slope is in the middle of this range on a logarithmic scale. Large-scale driving from galactic and extragalactic processes, including disk self-gravity, spiral waves and bars, presumably cause the low-k structure in what is effectively a two-dimensional geometry. Small-scale driving from stellar processes and shocks cause the high-k structure in a 3D geometry. This transition in dimensionality corresponds to the observed change in power spectrum slope. A companion paper models the observed power-law with a self-gravitating hydrodynamics simulation of a galaxy like the LMC.

Accepted for publication in The Astrophysical Journal Letters Available from arXiv:1006.2080

Sub-millimeter to centimeter excess emission from the Magellanic Clouds. I. Global spectral energy distribution

F.P. Israel^{1}, W.F. Wall^{2}, D. Raban^{1}, W.T. Reach^{3}, C. Bot^{4}, J.B.R. Oonk^{1}, N. Ysard^{5} and J.P.Bernard ^{6}

1Sterrewacht Leiden, The Netherlands 2INAOE, México 3California Institute of Technology, USA 4Observatoire Astronomique de Strasbourg, France 5University of Helsinki, Finland 6Université de Toulouse, France

In order to reconstruct the global SEDs of the Magellanic Clouds over eight decades in spectral range, we combined literature flux densities representing the entire LMC and SMC respectively, and complemented these with maps extracted from the WMAP and COBE databases covering the missing the 23–90 GHz (13–3.2 mm) and the poorly sampled 1.25–250 THz (240–1.25 μm). We have discovered a pronounced excess of emission from both Magellanic Clouds, but especially the SMC, at millimeter and sub-millimeter wavelengths. We also determined accurate thermal radio fluxes and very low global extinctions for both LMC and SMC. Possible explanations are briefly considered but as long as the nature of the excess emission is unknown, the total dust masses and gas-to-dust ratios of the Magellanic Clouds cannot reliably be determined.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:1006.2232

Far ultraviolet diffuse emission from the Large Magellanic Cloud

Ananta C. Pradhan^{1}, Amit Pathak^{1} and Jayant Murthy^{1}

1Indian Institute of Astrophysics, India

We present the first observations of diffuse radiation in the far ultraviolet (1000–1150 Å) from the Large Magellanic Cloud based on observations made with the Far Ultraviolet Spectroscopic Explorer. The fraction of the total radiation in the field emitted as diffuse radiation is typically 5–20% with a high of 45% near N 70 where there are few exciting stars, indicating that much of the emission is not due to nearby stars. Much less light is scattered in the far ultraviolet than at longer wavelengths with the stellar radiation going into heating the interstellar dust.

Accepted for publication in ApJ Letters Available from arXiv:1006.3912 and from http://www.iiap.res.in/people/personnel/acp/Research.html

Yellow supergiants in the Small Magellanic Cloud (SMC): Putting current evolutionary theory to the test

Kathryn F. Neugent^{1}, Philip Massey^{1}, Brian Skiff^{1}, Maria R. Drout^{2}, Georges Meynet^{3} and Knut A.G. Olsen^{4}

1Lowell Observatory, USA 2Department of Physics & Astronomy, University of Iowa, USA 3Geneva University, Geneva Observatory, Switzerland 4National Optical Astronomy Observatory, USA

The yellow supergiant content of nearby galaxies provides a critical test of massive star evolutionary theory. While these stars are the brightest in a galaxy, they are difficult to identify because a large number of foreground Milky Way stars have similar colors and magnitudes. We previously conducted a census of yellow supergiants within M 31 and found that the evolutionary tracks predict a yellow supergiant duration an order of magnitude longer than we observed. Here we turn our attention to the SMC, where the metallicity is 10× lower than that of M 31, which is important as metallicity strongly affects massive star evolution. The SMC’s large radial velocity (∼160 km s^{−1}) allows us to separate members from foreground stars. Observations of ∼500 candidates yielded 176 near-certain SMC supergiants, 16 possible SMC supergiants, along with 306 foreground stars and provide good relative numbers of yellow supergiants down to 12 M_{⊙}. Of the 176 near-certain SMC supergiants, the kinematics predicted by the Besançon model of the Milky Way suggest a foreground contamination of ≤4%. After placing the SMC supergiants on the H–R diagram and comparing our results to the Geneva evolutionary tracks, we find results similar to those of the M 31 study: while the locations of the stars on the H–R diagram match the locations of evolutionary tracks well, the models over-predict the yellow supergiant lifetime by a factor of ten. Uncertainties about the mass-loss rates on the main-sequence thus cannot be the primary problem with the models.

Accepted for publication in ApJ Available from arXiv:1006.5742

Exploring the origin and fate of the Magellanic Stream with UV and optical absorption

Andrew J. Fox^{1}, Bart P. Wakker^{2}, Jonathan V. Smoker^{1}, Philipp Richter^{3}, Blair D. Savage^{2} and Kenneth R. Sembach^{4}

1ESO, Chile 2UW-Madison, USA 3University of Potsdam, Germany 4STScI, USA

We present an analysis of ionization and metal enrichment in the Magellanic Stream (MS), the nearest gaseous tidal stream, using HST/STIS and FUSE ultraviolet spectroscopy of two background AGN, NGC 7469 and Mrk 335. For NGC 7469, we include optical spectroscopy from VLT/UVES. In both sightlines the MS is detected in lowion and high-ion absorption. Toward NGC 7469, we measure a MS oxygen abundance [O/H]_{MS} = [O i/H i] = −1.00 ± 0.05(stat)±0.08(syst), supporting the view that the Stream originates in the SMC rather than the LMC. We use cloudy to model the low-ion phase of the Stream as a photoionized plasma using the observed Si iii/Si ii and C iii/C ii ratios. Toward Mrk 335 this yields an ionization parameter log U between −3.45 and −3.15 and a gas density log(n_{H}/cm^{−3}) between −2.51 and −2.21. Toward NGC 7469 we derive sub-solar abundance ratios for [Si/O], [Fe/O], and [Al/O], indicating the presence of dust in the MS. The high-ion column densities are too large to be explained by photoionization, but also cannot be explained by a single-temperature collisional-ionization model (equilibrium or non-equilibrium). This suggests the high-ion plasma is multi-phase. Summing over the low-ion and high-ion phases, we derive conservative lower limits on the ratio N(total H ii)/N(H i) of > 19 toward NGC 7469 and > 330 toward Mrk 335, showing that along these two directions the vast majority of the Stream has been ionized. The presence of warm-hot plasma together with the small-scale structure observed at 21 cm provides evidence for an evaporative interaction with the hot Galactic corona. This scenario, predicted by hydrodynamical simulations, suggests that the fate of the MS will be to replenish the Galactic corona with new plasma, rather than to bring neutral fuel to the disk.

Accepted for publication in ApJ Available from arXiv:1006.0974

HERschel Inventory of The Agents of Galaxy Evolution (HERITAGE):

the Large Magellanic Cloud dust

M. Meixner^{1,2,3}, F. Galliano^{3}, S. Hony^{3}, J. Roman-Duval^{1}, T. Robitaille^{2}, P. Panuzzo^{3}, M. Sauvage^{3}, K. Gordon^{1}, C. Engelbracht^{4}, K. Misselt^{4}, K. Okumura^{3}, T. Beck^{1}, J.P. Bernard^{5}, A. Bolatto^{6}, C. Bot^{7}, M. Boyer^{1}, S. Bracker^{8},

L.R. Carlson^{9}, G.C. Clayton^{10}, C.-H. R. Chen^{11}, E. Churchwell^{8}, Y. Fukui^{12}, M Galametz^{3}, J.L. Hora^{2}, A. Hughes^{13}, R. Indebetouw^{11}, F.P. Israel^{14}, A. Kawamura^{12}, F. Kemper^{15}, S. Kim^{16}, E. Kwon^{16}, B. Lawton^{1}, A. Li^{17}, K.S. Long^{1}, M. Marengo^{18}, S.C. Madden^{3}, M. Matsuura^{19,20}, J.M.Oliveira^{21}, T. Onishi^{22}, M. Otsuka^{1}, D. Paradis^{23}, A. Poglitsch^{24}, D. Riebel^{9}, W.T. Reach^{23,25}, M. Rubio^{26}, B. Sargent^{1}, M. Sewiło^{1}, J.D. Simon^{27}, R. Skibba^{4}, L.J. Smith^{1}, S. Srinivasan^{28}, A.G.G.M. Tielens^{12}, J.Th. van Loon^{21}, B. Whitney^{29} and P.M. Woods^{15}

1Space Telescope Science Institute, USA 2Harvard-Smithsonian Center for Astrophysics, USA 3CEA, Laboratoire AIM, France 4Steward Observatory, University of Arizona, USA 5Centre d’Étude Spatiale des Rayonnements, France 6Department of Astronomy, Lab for Millimeter-wave Astronomy, University of Maryland, USA 7Observatoire Astronomique de Strasbourg, France 8Department of Astronomy, University of Wisconsin, USA 9Johns Hopkins University, Dept. of Physics & Astronomy, USA 10Lousianna State Univeristy, USA 11Department of Astronomy, University of Virginia, USA 12Department of Astrophysics, Nagoya University, Japan 13Centre for Supercomputing and Astrophysics, Swinburne University, Australia 14Sterrrewacht Leiden, Leiden University, The Netherlands 15Jodrell Bank Centre for Astrophysics, University of Manchester, UK 16Astronomy & Space Science, Sejong University, South Korea 17Department of Physics and Astronomy, University of Missouri, USA 18Department of Physics and Astronomy, Iowa State University, USA 19Department of Physics and Astronomy, University College London, UK 20MSSL, UCL, Holmbury, UK 21School of Physical & Geographical Sciences, Keele University, UK 22Department of Physical Science, Osaka Prefecture University, Japan 23Spitzer Science Center, California Institute of Technology, USA 24Max-Planck-Institut für extraterrestrische Physik, Germany 25Stratospheric Observatory for Infrared Astronomy, USA 26Departamento de Astronomía, Universidad de Chile, Chile 27Observatories of the Carnegie Institution of Washington, USA 28Institut d’Astrophysique de Paris, France 29Space Science Institute, USA

The HERschel Inventory of The Agents of Galaxy Evolution (HERITAGE) of the Magellanic Clouds will use dust emission to investigate the life cycle of matter in both the Large and Small Magellanic Clouds (LMC and SMC). Using the Herschel Space Observatory’s PACS and SPIRE photometry cameras, we imaged a 2^{∘} ×8^{∘} strip through the LMC, at a position angle of ∼ 22.5^{∘} as part of the science demonstration phase of the Herschel mission. We present the data in all 5 Herschel bands: PACS 100 and 160 μm and SPIRE 250, 350 and 500 μm. We present two dust models that both adequately fit the spectral energy distribution for the entire strip and both reveal that the SPIRE 500 μm emission is in excess of the models by ∼6 to 17%. The SPIRE emission follows the distribution of the dust mass, which is derived from the model. The PAH-to-dust mass (f_{PAH}) image of the strip reveals a possible enhancement in the LMC bar in agreement with previous work. We compare the gas mass distribution derived from the H i 21 cm and CO J = 1–0 line emission maps to the dust mass map from the models and derive gas-to-dust mass ratios (GDRs). The dust model, which uses the standard graphite and silicate optical properties for Galactic dust, has a very low GDR = 65_{−18}^{+15} making it an unrealistic dust model for the LMC. Our second dust model, which uses amorphous carbon instead of graphite, has a flatter emissivity index in the submillimeter and results in a GDR = 287_{−42}^{+25} that is more consistent with a GDR inferred from extinction. Accepted for publication in Astronomy & Astrophysics Available from arXiv:1006.0985

The pulsation of AGB stars in the Magellanic Cloud clusters NGC 1978 and NGC 419

D. Kamath^{1}, P.R. Wood^{1}, I. Soszyński^{2} and T. Lebzelter^{3}

1Research School of Astronomy & Astrophysics, Australian National University, Weston Creek, ACT 2611, Australia 2Warsaw University Observatory, Aleje Ujazdowskie 4, 00-478 Warszawa, Poland 3Institute of Astronomy, University of Vienna, Türkenschanzstraße 17, 1180 Vienna, Austria

The intermediate-age Magellanic Cloud clusters NGC 1978 and NGC 419 are each found to contain substantial numbers of pulsating AGB stars, both oxygen-rich and carbon-rich. Each cluster also contains two pulsating asymptotic giant branch (AGB) stars which are infrared sources with a large mass-loss rate. Pulsation masses have been derived for the AGB variables, from the lowest luminosity O-rich variables to the most evolved infrared sources. It is found that the stars in NGC 1978 have a mass of 1.55 M_{⊙} early on the AGB while the NGC 419 stars have a mass of 1.87 M_{⊙} early on the AGB. These masses are in good agreement with those expected from the cluster ages determined by main-sequence turnoff fitting. Nonlinear pulsation models fitted to the highly evolved AGB stars show that a substantial amount of mass loss has occurred during the AGB evolution of these stars. An examination of the observed mass loss on the AGB, and the AGB tip luminosities, shows that in both clusters the mass loss rates computed from the formula of Vassiliadis & Wood (1993) reproduce the observations reasonably well. The mass loss rates computed from the formula of Blöcker (1995) terminate the AGB in both clusters at a luminosity which is much too low.

Accepted for publication in MNRAS Available from arXiv:1006.3121

The ^{13}Carbon footprint of B[e] supergiants

A. Liermann^{1,2}, M. Kraus^{3}, O. Schnurr^{4,5} and M. Borges Fernandes^{6}

1Universität Potsdam, Institut für Physik und Astronomie, Karl-Liebknecht-Str. 24–25, 14476 Potsdam, Germany 2Max-Planck-Institut für Radioastronomie, Auf dem Hügel 69, 53121 Bonn, Germany 3Astronomický ústav, Akademie věd České republiky, Fričova 298, 251 65 Ondřejov, Czech Republic 4University of Sheffield, Department of Physics and Astronomy, University of Sheffield, Sheffield S3 7RH, UK 5Astrophysikalisches Institut Potsdam, An der Sternwarte 16, 14482 Potsdam, Germany 6Observatório Nacional, Rua General José Cristino 77, 20921-400 São Cristovão, Rio de Janeiro, Brazil

We report on the first detection of ^{13}C enhancement in two B[e] supergiants in the Large Magellanic Cloud. Stellar evolution models predict the surface abundance in ^{13}C to strongly increase during main-sequence and post-main sequence evolution of massive stars. However, direct identification of chemically processed material on the surface of B[e] supergiants is hampered by their dense, disk-forming winds, hiding the stars. Recent theoretical computations predict the detectability of enhanced ^{13}C via the molecular emission in ^{13}CO arising in the circumstellar disks of B[e] supergiants. To test this potential method and to unambiguously identify a post-main sequence B[e]SG by its ^{13}CO emission, we have obtained high-quality K-band spectra of two known B[e] supergiants in the Large Magellanic Cloud, using the Very Large Telescope’s Spectrograph for INtegral Field Observation in the Near-Infrared (VLT/SINFONI). Both stars clearly show the ^{13}CO band emission, whose strength implies a strong enhancement of ^{13}C, in agreement with theoretical predictions. This first ever direct confirmation of the evolved nature of B[e] supergiants thus paves the way to the first identification of a Galactic B[e] supergiant.

Accepted for publication in MNRAS Letters

The constant inner-disk radius of LMC X-3: A basis for measuring black hole spin

James F. Steiner^{1}, Jeffrey E. McClintock^{1}, Ronald A. Remillard^{2}, Lijun Gou^{1}, Shin’ya Yamada^{3} and Ramesh Narayan^{1}

1Harvard-Smithsonian CfA, 60 Garden St., Cambridge, MA 02138, USA 2MIT Kavli Institute for Astrophysics and Space Research, 70 Vassar Street, Cambridge, MA 02139, USA 3Department of Physics, University of Tokyo 7-3-1, Hongo, Bunkyo-ku, Tokyo 113-0033, Japan

The black-hole binary system LMC X-3 has been observed by virtually every X-ray mission since the inception of X-ray astronomy. Among the persistent sources, LMC X-3 is uniquely both habitually soft and highly variable. Using a fully relativistic accretion-disk model, we analyze hundreds of spectra collected during eight X-ray missions that span 26 years. For a selected sample of 391 RXTE spectra we find that to within ≈ 2 per cent the inner radius of the accretion disk is constant over time and unaffected by source variability. Even considering an ensemble of eight X-ray missions, we find consistent values of the radius to within ≈ 4–6 per cent. Our results provide strong evidence for the existence of a fixed inner-disk radius. The only reasonable inference is that this radius is closely associated with the general relativistic innermost stable circular orbit (ISCO). Our findings establish a firm foundation for the measurement of black hole spin.

Accepted for publication in ApJL Available from arXiv:1006.5729

Radio emission from dark matter annihilation in the Large Magellanic Cloud

Beatriz B. Siffert^{1}, Angelo Limone^{1}, Enrico Borriello^{1,2}, Giuseppe Longo^{1,3,4} and Gennaro Miele^{1,2}

1Dipartimento di Scienze Fisiche, Università di Napoli Federico II, Via Cinthia, 9, I-80126, Naples, Italy 2INFN — Sezione di Napoli, Via Cinthia, 9, I-80126, Naples, Italy 3INAF-OACN, Via Moiariello, 16, I-80128, Naples, Italy 4Department of Astronomy, California Institute of Technology, 1200 East California Blvd, Pasadena, CA 91125, USA

The Large Magellanic Cloud, at only 50 kpc away from us and known to be dark matter dominated, is clearly an interesting place where to search for dark matter annihilation signals. In this paper, we estimate the synchrotron emission due to WIMP annihilation in the halo of the LMC at two radio frequencies, 1.4 and 4.8 GHz, and compare it to the observed emission, in order to impose constraints in the WIMP mass vs. annihilation cross section plane. We use available Faraday rotation data from background sources to estimate the magnitude of the magnetic field in different regions of the LMC’s disc, where we calculate the radio signal due to dark matter annihilation. We account for the e^{+}e^{−} energy losses due to synchrotron, Inverse Compton Scattering and bremsstrahlung, using the observed hydrogen and dust temperature distribution on the LMC to estimate their efficiency. The extensive use of observations, allied with conservative choices adopted in all the steps of the calculation, allow us to obtain very realistic constraints.

Submitted to MNRAS Available from arXiv:1006.5325

Interstellar absorptions towards the LMC: Small-scale density variations in Milky Way disc gas

Soroush Nasoudi-Shoar^{1}, Philipp Richter^{2}, Klaas S. de Boer^{1} and Bart P. Wakker^{3}

1Argelander-Institut für Astronomie, Universität Bonn, Auf dem Hügel 71, 53121 Bonn, Germany 2Institut für Physik und astronomie, Universität Potsdam, Haus 28, Karl-Liebknecht-Str. 24/25, 14476 Golm, Germany 3Department of Astronomy, University of Wisconsin, 475 N Charter St., Madison, WI 53706

Observations show that the ISM contains sub-structure on scales less than 1 pc, detected in the form of spatial and temporal variations in column densities or optical depth. Despite the number of detections, the nature and ubiquity of the small-scale structure in the ISM is not yet fully understood. We use UV absorption data mainly from FUSE and partly from STIS of six LMC stars (Sk-67^{∘}111, LH 54-425, Sk-67^{∘}107, Sk-67^{∘}106, Sk-67^{∘}104, and Sk-67^{∘}101), all located within 5′ of each other, and analyse the physical properties of the Galactic disc gas in front of the LMC on sub-pc scales. We analyse absorption lines of a number of ions within the UV spectral range. Most importantly, interstellar molecular hydrogen, neutral oxygen, and fine-structure levels of neutral carbon have been used in order to study changes in the density and the physical properties of the Galactic disc gas over small angular scales. While most species do not show any significant variation in their column densities, we find an enhancement of almost 2 dex for H_{2} from Sk-67^{∘}111 to Sk-67^{∘}101, accompanied by only a small variation in the O i column density. Based on the formation–dissociation equilibrium, we trace these variations to the actual density variations in the molecular gas. On the smallest spatial scale of < 0.08 pc, between Sk-67^{∘}107 and LH 54-425, we find a gas density variation of a factor of 1.8. The line of sight towards LH 54-425 does not follow the relatively smooth change seen from Sk-67^{∘}101 to Sk-67^{∘}111, suggesting that sub-structure might exist on a smaller spatial scale than the linear extent of our sight-lines. Our observations suggest that the detected H_{2} in these six lines of sight is not necessarily physically connected, but that we are sampling molecular cloudlets with pathlengths < 0.1–1.8 pc and possibly different densities.

Accepted for publication in A&A Available from arXiv:1006.5450

The mid-infrared period-luminosity relations for the Small Magellanic Cloud Cepheids derived from Spitzer archival data

Chow-Choong Ngeow^{1} and Shashi M. Kanbur^{2}

1National Central University. Taiwan 2State University of New York at Oswego, USA

In this paper we derive the Spitzer IRAC band period–luminosity (P–L) relations for the Small Magellanic Cloud (SMC) Cepheids, by matching the Spitzer archival SAGE-SMC data with the OGLE-III SMC Cepheids. We find that the 3.6 μm and 4.5 μm band P–L relations can be better described using two P–L relations with a break period at log P = 0.4: this is consistent with similar results at optical wavelengths for SMC P–L relations. The 5.8 μm and 8.0 μm band P–L relations do not extend to sufficiently short periods to enable a similar detection of a slope change at log P = 0.4. The slopes of the SMC P–L relations, for log P > 0.4, are consistent with their LMC counterparts that were derived from a similar dataset. They are also in agreement with those obtained from a small sample of Galactic Cepheids with parallax measurements.

Accepted for publication in Astrophysical Journal Available from arXiv:1007.1499

A Galactic Center origin for HE 0437−5439, the hypervelocity star near the Large Magellanic Cloud

Warren R Brown^{1}, Jay Anderson^{2}, Oleg Y. Gnedin^{3}, Howard E. Bond^{2}, Margaret J. Geller^{1}, Scott J. Kenyon^{1} and Mario Livio^{2}

1Smithsonian Astrophysical Observatory, USA 2Space Telescope Science Institute, USA 3University of Michigan, USA

We use Hubble Space Telescope imaging to measure the absolute proper motion of the hypervelocity star (HVS) HE 0437−5439, a short-lived B star located in the direction of the Large Magellanic Cloud (LMC). We observe (μ_{α}, μ_{δ}) = (+0.53±0.25(stat)±0.33(sys), +0.09±0.21(stat)±0.48(sys)) mas yr^{−1}. The velocity vector points directly away from the center of the Milky Way; an origin from the center of the LMC is ruled out at the 3-σ level. The flight time of the HVS from the Milky Way exceeds its main-sequence lifetime, thus its stellar nature requires it to be a blue straggler. The large space velocity rules out a Galactic-disk ejection. Combining the HVS’s observed trajectory, stellar nature, and required initial velocity, we conclude that HE 0437−5439 was most likely a compact binary ejected by the Milky Way’s central black hole.

Published in ApJ Letters Available from arXiv:1007.3493

Period–luminosity relations of pulsating M giants in the solar neighbourhood and the Magellanic Clouds

V. Tabur^{1}, T.R. Bedding^{1}, L.L. Kiss^{1,2}, T. Giles^{1}, A. Derekas^{1,2,3} and T.T. Moon^{1}

1Sydney Institute for Astronomy (SIfA), School of Physics, The University of Sydney, NSW 2006, Australia 2Konkoly Observatory, Hungarian Academy of Sciences, P.O. Box 67, H-1525 Budapest, Hungary 3Magyary Fellow, Dept. of Astronomy, Eötvös University, Budapest, Hungary

We analyse the results of a 5.5-yr photometric campaign that monitored 247 southern, semiregular variables with relatively precise Hipparcos parallaxes to demonstrate an unambiguous detection of Red Giant Branch (RGB) pulsations in the solar neighbourhood. We show that Sequence A′ contains a mixture of AGB and RGB stars, as indicated by a temperature related shift at the TRGB. Large Magellanic Cloud (LMC) and Galactic sequences are compared in several ways to show that the P–L sequence zero-points have a negligible metallicity dependence. We describe a new method to determine absolute magnitudes from pulsation periods and calibrate the LMC distance modulus using Hipparcos parallaxes to find μ_{LMC} = 18.54 ± 0.03 mag. Several sources of systematic error are discussed to explain discrepancies between the MACHO and OGLE sequences in the LMC. We derive a relative distance modulus of the Small Magellanic Cloud (SMC) relative to the LMC of Δμ = 0.41 ± 0.02 mag. A comparison of other pulsation properties, including period–amplitude and luminosity–amplitude relations, confirms that RGB pulsation properties are consistent and universal, indicating that the RGB sequences are suitable as high-precision distance indicators. The M giants with the shortest periods bridge the gap between G and K giant solar-like oscillations and M-giant pulsation, revealing a smooth continuity as we ascend the giant branch.

Accepted for publication in MNRAS Available from arXiv:1007.2974

Cosmic-ray acceleration efficiency vs. temperature equilibration: the case of SNR 0509−67.5

E.A. Helder^{1}, D. Kosenko^{1} and J. Vink^{1}

1Astronomical Insitute Utrecht, Utrecht University, The Netherlands

We study the 0509−67.5 supernova remnant in the Large Magellanic Cloud with the VLT/FORS2 spectrograph. We detect a broad component in the Hα emission with a FWHM of 2680 ± 70 km s^{−1} and 3900 ± 800 km s^{−1} for the SouthWest (SW) and NorthEast (NE) shocks respectively. For the SW, the proton temperature appears to be too low for the shock velocity, which we attribute to a cosmic-ray pressure behind the shock front of at least 20% of the total pressure. For the NE, the post-shock proton temperature and the shock velocity are compatible, only if the plasma behind the shock front has a degree of thermal equilibrium of over 20%, which is at odds with current models for temperature equilibration behind fast shocks, which do not accelerate cosmic rays. If we assume the electron temperature to be less than 10% of the proton temperature, we find a post-shock cosmic-ray pressure of at least 7%.

Accepted for publication in The Astrophysical Journal Letters Available from arXiv:1007.3138

XMM-Newton X-ray spectra of the SNR 0509−67.5: data and models

D. Kosenko^{1}, J. Vink^{1}, S. Blinnikov^{2} and A. Rasmussen^{3}

1Astronomical Institute Utrecht, The Netherlands 2Institute for Theoretical and Experimental Physics, Russia 3Stanford Linear Accelerator Center, CA, USA

Aims: We report on X-ray observations of the supernova remnant 0509−67.5 in the Large Magellanic Cloud acquired by the XMM-Newton X-ray observatory. We use the imaging spectroscopy (EPIC) and Reflective Grating Spectrometer (RGS) data to investigate properties of the remnant and its environment. Methods: The X-ray spectra were analyzed with SPEX software package. In addition, we performed a numerical hydrodynamic simulation of the remnant. Results: The EPIC data show prominent Fe K line emission, but the deduced overall amount of iron in the shocked ejecta is low. The data also show that the remnant has an asymmetric ejecta structure: the bright southwest region of the remnant shows an overabundance of metals. The analysis of the RGS spectrum shows that the remnant has a high line velocity broadening of 5000 km s^{−1}. We developed a hydrodynamical model for the remnant with basic hydrodynamical and spectral parameters that are similar to those observed. Conclusions: The data analysis indicates that the reverse shock just recently reached iron layers of the ejecta. The brightness enhancement in the southwest region could be a sign of an asymmetric explosion or it could be the result of a density enhancement in the interstellar medium. We constructed numerical models that are in good agreement with the observations, with circumstellar density of 3 × 10^{−25} g cm^{−3}, age of ∼ 400 years, velocities of ∼ 5000 km s^{−1}, and an electron-to-ion temperature ratio of 10^{−2}.

Published in Astronomy and Astrophysics, 490, 223 (2008) Available from arXiv:0807.0579

The R 136 star cluster hosts several stars whose individual masses greatly exceed the accepted 150 M_{⊙} stellar mass limit

Paul Crowther^{1}, Olivier Schnurr^{1,2}, Raphael Hirschi^{3,4}, Nothasliza Yusof^{5}, Richard Parker^{1}, Simon Goodwin^{1} and Hassan Kassim^{5}

1Sheffield, UK 2AIP, Germany 3Keele, UK 4Tokyo, Japan 5Malaya, Malaysia

Spectroscopic analyses of H-rich WN5–6 stars within the young star clusters NGC 3603 and R 136 are presented, using archival HST and VLT spectroscopy, and high spatial resolution near-IR photometry. We derive high stellar temperatures for the WN stars in NGC 3603 (T_{⋆} ∼ 42 ± 2 kK) and R 136 (T_{⋆} ∼ 53 ± 3 kK) plus clumping-corrected dM/dt ∼ 2–5×10^{−5} M_{⊙} yr^{−1} which closely agree with theoretical predictions. These stars make a disproportionate contribution to the global budget of their host clusters. R 136 a1 alone supplies ∼ 7% of N(LyC) of the entire 30 Dor region. Comparisons with stellar models calculated for the main-sequence evolution of 85–500 M_{⊙} suggest ages of ∼ 1.5 Myr and M_{init} in the range 105–170 M_{⊙} for 3 systems in NGC 3603, plus 165–320 M_{⊙} for 4 stars in R 136. Our high stellar masses are supported by dynamical mass determinations for the components of NGC 3603 A1. We consider the predicted L_{X} of the R 136 stars if they were close, colliding wind binaries. R 136 c is consistent with a colliding wind binary system. However, short period, colliding wind systems are excluded for R 136 a WN stars if mass ratios are of order unity. Widely separated systems would have been expected to harden owing to early dynamical encounters with other massive stars in such a dense environment. From simulated star clusters, whose constituents are randomly sampled from the Kroupa IMF, both clusters are consistent with a tentative upper mass limit of ∼ 300 M_{⊙}. The Arches cluster is either too old, exhibits a deficiency of very massive stars, or more likely stellar masses have been underestimated — M_{init} for the most luminous stars in the Arches cluster approach 200 M_{⊙} according to contemporary stellar and photometric results. The potential for stars greatly exceeding 150 M_{⊙} within metal-poor galaxies suggests that such pair-instability SNe could occur within the local universe, as has been claimed for SN 2007bi.

Accepted for publication in MNRAS Available from arXiv:1007.3284 and from http://pacrowther.staff.shef.ac.uk/R136.pdf

Infrared period–luminosity relations of evolved variable stars in the Large Magellanic Cloud

David Riebel^{1}, Margaret Meixner^{2}, Oliver Fraser^{3}, Sundar Srinivasan^{4}, Kem Cook^{5} and Uma Vijh^{6}

1Johns Hopkins University, USA 2Space Science Telescope Institute, USA 3University of Washington, USA 4Institut d’Astrophyisque de Paris, France 5Lawrence Livermore National Laboratory, USA 6University of Toledo, USA

We combine variability information from the MAssive Compact Halo Objects (MACHO) survey of the Large Magellanic Cloud (LMC) with infrared photometry from the Spitzer Space Telescope Surveying the Agents of a Galaxy’s Evolution (SAGE) survey to create a dataset of ∼ 30 000 variable red sources. We photometrically classify these sources as being on the first ascent of the Red Giant Branch (RGB), or as being in one of three stages along the Asymptotic Giant Branch (AGB): oxygen-rich, carbon-rich, or highly reddened with indeterminate chemistry (“extreme” AGB candidates). We present linear period–luminosity relationships for these sources using 8 separate infrared bands (J, H, K, 3.6, 4.5, 5.8, 8.0, and 24 μm) as proxies for the luminosity. We find that the wavelength dependence of the slope of the period–luminosity relationship is different for different photometrically determined classes of AGB stars. Stars photometrically classified as O-rich show the least variation of slope with wavelength, while dust enshrouded extreme AGB stars show a pronounced trend toward steeper slopes with increasing wavelength. We find that O-rich AGB stars pulsating in the fundamental mode obey a period–magnitude relation with a slope of −3.41 ± 0.04 when magnitude is measured in the 3.6-μm band, in contrast to C-rich AGB stars, which obey a relation of slope −3.77 ± 0.05.

Accepted for publication in The Astrophysical Journal Available from arXiv:1007.5029

Star formation history and X-ray binary populations: The case of the Small Magellanic Cloud

Vallia Antoniou^{1,2,3}, Andreas Zezas^{1,2,4}, Despina Hatzidimitriou^{4,5} and Vicky Kalogera^{6}

1Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, USA 2Physics Department, University of Crete, P.O. Box 2208, GR-710 03, Heraklion, Crete, Greece 3Department of Physics and Astronomy, Iowa State University, 12 Physics Hall, Ames, IA 50011, USA 4IESL/Foundation for Research and Technology-Hellas, P.O. Box 1527, GR-711 10 Heraklion, Crete, Greece 5Section of Astrophysics, Astronomy, and Mechanics, Department of Physics, University of Athens, Panepistimiopolis, GR-157 84 Zografos, Athens, Greece 6Department of Physics and Astronomy, Northwestern University, 2145 Sheridan Road, Evanston, IL 60208, USA

Using Chandra, XMM-Newton, and optical photometric catalogs we study the young X-ray binary (XRB) populations of the Small Magellanic Cloud. We find that the Be/X-ray binaries (Be-XRBs) are observed in regions with star formation rate bursts ∼ 25–60 Myr ago. The similarity of this age with the age of maximum occurrence of the Be phenomenon (∼ 40 Myr) indicates that the presence of a circumstellar decretion disk plays a significant role in the number of observed XRBs in the 10–100 Myr age range. We also find that regions with strong but more recent star formation (e.g., the Wing) are deficient in Be-XRBs. By correlating the number of observed Be-XRBs with the formation rate of their parent populations, we measure a Be-XRB production rate of ∼ 1 system per 3 × 10^{−3} M_{⊙} yr^{−1}. Finally, we use the strong localization of the Be-XRB systems in order to set limits on the kicks imparted on the neutron star during the supernova explosion.

Published in ApJ, 716, L140 (2010) Available from arXiv:1006.2381

Conference Paper

Heated dust around the LMC Wolf–Rayet system HD 36402 (BAT 99-38)

P.M. Williams^{1}

1Institute for Astronomy, Royal Observatory Edinburgh, Scotland

Infrared photometry of the probable triple WC4 (+O?) +O8I: system HD 36402 (BAT 99-38) in the LMC shows emission characteristic of heated dust. Although HD 36402 is close to two luminous YSOs, it is possible to distinguish its emission at wavelengths less than 10 μm. Simple modelling indicates a dust temperature near 800 K and mass of about 1.5 × 10^{−7} M_{⊙} amorphous carbon grains. The dust emission appears to be variable. It is apparent that Wolf–Rayet dust formation occurs also in metal-poor environments.

Poster contribution, published in 39^{th} Liège International Astrophysical Colloquium Available from arXiv:1007.3645

Review Paper

Spectropolarimetry of Wolf–Rayet stars in the Magellanic Clouds: Constraining the progenitors of gamma-ray bursts

Jorick S. Vink^{1}

1Armagh Observatory, Northern Ireland

Wolf–Rayet stars have been identified as objects in their final phase of massive star evolution. It has been suggested that Wolf–Rayet stars are the progenitors of long-duration gamma-ray bursts in low-metallicity environments. However, this deduction has yet to be proven. Here we report on our initial results from a VLT/FORS linear spectropolarimetry survey of Wolf–Rayet stars in the Magellanic Clouds, which is intended to constrain the physical criteria — such as weaker stellar winds, rapid rotation, and associated asymmetry — of the collapsar model. Finally, we provide an outlook for polarisation studies with an extremely large telescope.

Published in Msngr 140, 46 (2010) Available from arXiv:1007.4336 and from http://www.eso.org/sci/publications/messenger/archive/no.140-jun10/messenger-no140-46-49.pdf

Job Advert

University of Hertfordshire University Research Fellowships

University of Hertfordshire School of Physics, Astronomy & Mathematics and Centre for Astrophysics Research

Applications are invited for a number of University Research Fellows in the Centre for Astrophysics Research, within the School of Physics, Astronomy & Mathematics. Research interests should closely match those within the Centre, which cover a wide range of Galactic and Extragalactic projects. Four posts are available, each for 2 years in the first instance, and can start from September 1, 2010.

Candidates should have a PhD, or expect to be awarded a PhD before the start date. The appointment will be on a salary ranging from £25,000 to £28,000, but might be higher for exceptional candidates.

Informal enquiries to Professor James Hough, Director of Astronomy Research, Centre for Astrophysics Research (telephone: +44 1707 284500, fax: +44 1707 284514; e-mail: j.h.hough@herts.ac.uk. Also see http://www.star.herts.ac.uk for additional information on research within the Centre.

Further details of the above post and application forms can be obtained on-line at http://recruitment.herts.ac.uk/recruit/ or from the Personnel Department at the University of Hertfordshire, College Lane, Hatfield, Hertfordshire AL10 9AB, England, or telephone 01707 284802, quoting reference number 000664. Applications (which should be accompanied by a full CV and statement of research interests) should be returned to Professor James Hough or to the Personnel Department at the above address.

Applicants should arrange for 2 references to be sent to Professor James Hough or to Personnel by the closing date of August 6, 2010.

The University is an equal opportunity employer.

Additional note for this Newsletter:

Applications are encouraged to work on the VISTA survey of the Magellanic system (VMC) survey.

For more information about the survey science contact the PI at M.Cioni@herts.ac.uk

Announcement

Magellanic Clouds meeting

Dear colleagues, I would like to inform you that a meeting dedicated to the observational/theoretical research of the Magellanic Clouds is scheduled as a part of the conference ”Zooming in: The Cosmos at High Resolution, Bonn, Germany, 13–17 Sept. 2010”. For detailed information and registration, please visit the conference website http://www.astro.uni-bonn.de/AG201/index.html Information about the Magellanic Clouds meeting is available in http://homepage.univie.ac.at/adam.ruzicka/AG2010magellanic Best wishes,

Adam Růžička Institut für Astronomie, Universität Wien, Austria adam.ruzicka@univie.ac.at See also http://www.astro.uni-bonn.de/AG2010/splinters.html